Determine the value of K for the balanced reaction, 2 NOCl(g) ↔ N2(g) + O2(g) + Cl2(g) at 291 K given the following information,

2 N2(g) + 2 O2(g) ↔ 4 NO(g); K = 5.397×10-61

NO + 1/2 Cl2(g) ↔ NOCl(g); K = 2.432×104
Answer #1

Solukoy - 2N2 g + 2020, = > k= [noze 4Nog; = 5.39781367 =) (nole (1122? [02]? 5.392x100 Noot Ź cl g = Noug ; K= 2-432710 7 K=a pa ][0:1][4) K 0 Takip square root af eq6) on both sids en meng = 7.301810-9 Taky reciprocal on both sids => [m2][2] = 1361